OpenStudy (anonymous):

Prove the identity. a^8-b^8=(a-b)(a+b)(a^2+b^2)(a^4+b^4)

OpenStudy (callisto):

Start from the right side RS = (a-b)(a+b)(a^2+b^2)(a^4+b^4) First, use difference of two squares to expand (a-b)(a+b), ie. (a-b)(a+b) = (a^2 - b^2) So, you get RS = (a-b)(a+b)(a^2+b^2)(a^4+b^4) = (a^2 - b^2))(a^2+b^2)(a^4+b^4) Use the same trick to expand it. Can you do it now?

OpenStudy (anonymous):

Yes Thanks!

OpenStudy (callisto):

Welcome :)

OpenStudy (callisto):

You can also start from the left though a^8-b^8 = (a^4)^2 - (b^4)^2 = (a^4 - b^4)(a^4 + b^4) = [ (a^2)^2 - (b^2)^2 ] (a^4 + b^4) Then, difference of two squares again...